A man who can walk 4 miles per hour takes a walk into the country on a holiday. He is away for 6 hours. How far does he go if he return by bicycle at the average rate of 12 miles an hour?
Answer by mananth(12270)

walk 4mph
cycle 12 mph
let distance be x miles
..
time taken walking = x/4
..
time taken riding back = x/12
x/12+x/4 = 6
(x+3x)/12=6
4x=72
/4
x=18 miles.
...
m.ananth@hotmail.ca